# Planning Shorter Paths in Graphs of Convex Sets by Undistorting Parametrized Configuration Spaces

### v1

@@ Abstract
Optimization based motion planning provides a useful modeling framework through various costs and constraints. Using Graph of Convex Sets (GCS) for trajectory optimization gives guarantees of feasibility and optimality by representing configuration space as the finite union of convex sets. Nonlinear parametrizations can be used to extend this technique to handle cases such as kinematic loops, but this distorts distances, such that solving with convex objectives will yield paths that are suboptimal in the original space. We present a method to extend GCS to nonconvex objectives, allowing us to “undistort” the optimization landscape while maintaining feasibility guarantees. We demonstrate our method’s efficacy on three different robotic planning domains: a bimanual robot moving an object with both arms, the set of 3D rotations using Euler angles, and a rational parametrization of kinematics that enables certifying regions as collision free. Across the board, our method significantly improves path length and trajectory duration with only a minimal increase in runtime.
@@ I Introduction
Reliable motion planning is essential to developing and deploying robotic manipulation systems. Such systems need to produce efficient paths while obeying various constraints. Optimization-based motion planning, which tries to minimize an objective function while satisfying constraints, offers a powerful paradigm to solve this problem. The decision variables describe the robot’s trajectory, the objective allows for choosing desired qualities in the solution, and constraints on these decision variables define obstacle avoidance, dynamic limits, and other interesting task-specific constraints such as coordinating arms in a bimanual system. However, the power of these techniques is tempered by the need to carefully formulate the optimization problems for reliability.
@@
A manipulator’s configuration space is often inherently nonconvex, and nonconvex trajectory optimization usually cannot guarantee optimality (due to local minima), or even feasibility. These guarantees are key to efficient and robust systems that can be used for repetitive motions in safety critical settings. Graphs of Convex Sets (GCS) [1, 2] encapsulates the nonconvexities from obstacle avoidance as discrete decisions, casting it as a convex optimization problem. More concretely, it takes the inner approximation of planning or configuration space represented as a series of intersecting collision free convex subsets. Then, it can optimize over a graph (where vertices are a convex c-space set and edges represent intersections between these sets) for discrete paths through these sets and simultaneously optimize for the shortest or fastest continuous path within each set. Many relevant properties of the trajectory and its derivatives can be transcribed into convex costs and constraints [2].
@@
Figure 1: Experiments include constrained bimanual motion planning between shelves (top) and certifiable 7DoF KUKA iiwa trjaectories between bins (below). The red path is the original result, and the blue path is our improved result.
When the configuration space does not admit finite Euclidean inner approximations, recent work has reformulated the problem through parametrization to use GCS. For example, Cohn et al. [3] tackle planning with an equality constraint between end effectors of two manipulators using inverse kinematics (IK). This parametrization satisfies the aforementioned equality constraint by construction [3]. Another useful application of parametrizations is the Euler angles representation of 3D rotations, which enables planning over this space with GCS [4, §2.7.5]. Certified C-IRIS [5] uses a rational parametrization of robot kinematics to formulate an optimization problem to provide collision free certificates in presence of task space obstacles.
@@
Unfortunately, these parametrizations are non-isometric; they distort space, so the shortest path in the parametrized space may not be the shortest path in the original configuration space. This leads to suboptimal results when the convex objective used (path length costs) in the parametrized space is a weak approximation for the true objective. Allowing GCS to handle nonconvex objectives, such as the true objective in the parametrized space, gives more modeling freedom, widening the breadth of problems we can tackle while maintaining guarantees.
@@
This work considers GCS problems with nonconvex objectives, while maintaining convex constraints. Specifically, we use Projected Gradient Descent (PGD) to optimize such nonconvex objectives in a GCS approach. Using a gradient based solver guarantees local optimality around the initial guess if the objective is Lipschitz-continuous. Since we keep constraints convex, a small convex program can always project infeasible solutions back to feasibility. Thus by exploiting structured nonconvexity, our solver improves optimality of solutions while maintaining feasibility guarantees.
@@
For each of the three parametrizations mentioned earlier (bimanual IK, Euler angles, rational kinematics), we formulate and test a nonconvex objective against a convex surrogate in the GCS optimization formulation. This nonconvex optimization is treated as a postprocessing step, improving upon the best solution from GCS. Our method offers significant quantitative and qualitative improvements to motion plans across multiple experiments: path lengths and trajectory times shorten and visual artifacts of planning in the distorted parametrized space are undone. Expanding beyond the regime of undoing distortions, we also optimize over a general nonconvex cost, spatial curvature, to further improve the results from the bimanual experiment.
@@
In the rest of this paper, we first review related work on nonconvex trajectory optimization and give necessary background on GCS and existing efforts to add nonconvexity. Then, we describe our methodology, relevant implementation details, and experimental set-ups. Finally, we present our experimental results, and conclude with a brief discussion of the limitations of our work and potential directions for future research.
@@ II-A Graphs of Convex Sets
Graph of Convex Sets (GCS) presents a new strategy for solving shortest path problems with continuous and discrete decisions. Formally, a GCS is a graph, where each vertex $v$ has an associated continuous variable $x_{v}$ within a convex set $X_{v}$, and each edge $(u,v)$ is a convex function of $x_{u}$ and $x_{v}$. Finding the shortest path $\mathcal{P}$ through this graph can then be formulated as a Mixed Integer Convex Problem (MICP) with $\mathcal{P}$ and $x_{v}$ as decision variables [1, §5].
In motion planning, the intuitive application of GCS is to plan through overlapping convex sets and then optimize trajectories within the sets. These sets constitute the vertices of $\mathcal{P}$, and their union is an inner approximation of our planning space. $x_{v}$ or the continuous variable within each set then is the continuous trajectory through it. Like the paper introducing GCS for Trajectory Optimization paper (GCSTrajOpt) [2] we define $x_{v}$ as the control points of a Bézier curve to parametrize the continuous trajectory. This choice admits convex path continuity and differentiability constraints, and collision-avoidance of the whole trajectory is guaranteed [2, p.9]. To optimize for the shortest path, GcsTrajOpt minimizes the distance between adjacent control points as a proxy for minimizing the length of the curve [1]. We use the same objective for the convex relaxation and any convex optimization we do.
The above discussion focuses only on the path, but in GcsTrajOpt, $x_{v}$ has an additional parameter: a time-scaling variable, $h$. We do not use this additional parameter to avoid nonconvex acceleration constraints that use this time parametrization variable to ensure our trajectories obey physical limits. Instead we use TOPP (Time Optimal Path Parametrization) [17] to generate timed trajectories from our planned spatial paths. As any differentiable path can be navigated under acceleration constraints by slowing down, using TOPP helps maintain feasibility guarantees by keeping nonconvexity out of our constraints. So, for a collision free convex set $Q_{i}$, our vertices are of dimension $Q_{i}^{d+1}$ given Bézier curves of degree $d$ with $d+1$ control points.
@@ II-B1 IK and Constrained Bimanual Planning
Constrained bimanual manipulation refers to two robot arms that move with a fixed transform between their end effectors. This nonlinear equality constraint in task space prevents using GCS to plan in the full $\mathbb{R}^{14}$ configuration space. Cohn et al. [3] use inverse kinematics to determine the joint angles of a subordinate robot given the end effector position of the leading arm. This parametrization collapses the $\mathbb{R}^{14}$ full joint space over both arms into an $\mathbb{R}^{8}$ space formed by the leading arm’s joints and a redundancy factor (required to generate consistent joint angles for the subordinate arm). Thus, no nonlinear equality constraints are required. However, trying to minimize path length in the $\mathbb{R}^{8}$ effectively minimizes the path length for the leading arm while ignoring the subordinate arm. As a result, paths are visibly imbalanced and favor the leading arm.
@@ II-B3 Rational Kinematics to Certify Collision Free
The forward kinematic mapping (needed for checking if a configuration is collision-free) is a trigonometric polynomial. Amice et al. [5] transform this nonconvex relationship into a multi-linear polynomial, using the tangent half-angle substitution $s=\tan\frac{\theta}{2}$, further implying
@@
for $\theta\in(-\pi,\pi)$. They then use this change of coordinates to formulate Semi-Definite Programs (SDPs) that can certify non-collision with task space obstacles for regions in the robot’s rational configuration space. This certification can be done for individual convex sets [5], or even an entire trajectory [18].
Figure 2: A stereographic projection about $N$ will project the bottom of the black rectangle as being smaller than the top. A planner operating in the post-projection (parametrized) space and optimizing for short distances would favor the bottom even though both the sides are equal in the original setting. Image generated using code from [19].
@@ II-C Nonconvexity and GCS
There is precedent for working with nonconvexity in GCS or similar optimization frameworks. The original GcsTrajOpt preprint [20] suggested using convex approximations to incorporate nonconvex objectives and constraints. In line with this suggestion, existing GCS works using the parametrizations from Subsection II-B use a convex surrogate objective to approximate the optimal solution. While this approach preserves convexity, the approximations are inherently heuristic and must be hand-designed. Moreover, optimizing for an approximation means that the optimality of the solution is bounded by the quality of the approximation. The nuances of the approximations can also lead to systematic pathologies, such as the imbalance between arms in the bimanual planning domain.
Another approach is to work with local convex approximations of the nonconvexity. Clark and Xie [21] suggest approximating the nonconvex costs using piecewise-linear approximations and creating smaller sets within which the objective is convex. This approach maintains convexity, but may scale poorly when dealing with complex objectives and finer approximations. Using a mix of biconvex alternation and local convex approximation, Fast Path Planning [22] handles a bilinear in a similar set-up as GCS. The nonconvexity in this problem is contained to the constraints and is handled using alternation. The nonconvexity being a bilinear nonconvexity is key to enabling this method. Our work aims to enable a broader class of nonconvexity in the objective functions.
@@ III-A Nonlinear Changes of Coordinates
Each of the parametrizations discussed in Subsection II-B distort the robot’s configuration space by introducing a nonlinear change of coordinates. More formally, each domain has a smooth (nonlinear) transformation $\alpha:Q\rightarrow C$ that maps $x$ from the more useful parametrized space $Q$ to a point $\alpha(x)$ in the original configuration space $C$. Each of the works used GCS to solve for a trajectory in $Q$, which then is remapped to $C$ using $\alpha$ to get an actual robot trajectory. However, since $\alpha$ is a nonlinear transformation, the path with minimum length in $Q$ is not guaranteed to be the path of minimum length in $C$.
The key limitation is that the convex path length cost in $Q$ can be arbitrarily far from the true objective: minimizing distance in $C$. Using $\alpha$ enables changing coordinates back to the original space $C$ and defining this true (now nonconvex) objective, but the sets and constraints stay convex in the parametrized space $Q$.
For the constrained bimanual case, $\alpha:\mathbb{R}^{8}\rightarrow\mathbb{R}^{14}$ is defined as the nonlinear analytic IK function with an original configuration space of both arms’ joints ($\mathbb{R}^{14}$) and a parametrized space of one arm’s joints and the self-motion of the other arm ($\mathbb{R}^{8}$). For planning over $\operatorname{SO}(3)$ with Euler angles, $\alpha:\mathbb{R}^{3}\rightarrow\mathbb{R}^{4}$ is the standard conversion from Euler angles to quaternions. For planning in the rational parametrization of kinematics, $\alpha$ is defined as $\theta=2\tan^{-1}s$.
@@ III-B Formulating the Optimization
The nonconvex objective using $\alpha$ still needs to be expressed in terms of our decision variables $x_{v}$, the control points of the Bézier curve in $Q$. We cannot directly apply $\alpha$ on $x_{v}$ to define a distance objective as the remapped control points from $Q$ do not define a same Bézier curve in $C$. However, any points along the Bézier curve in $Q$ will still be along the same path in $C$, and any point along the Bézier curve is a convex combination of its control points. Therefore, a piecewise-linear approximation of the curve in $Q$ can be mapped to a piecewise-linear approximation in $C$ using $\alpha$.
The representative cost can then be the length of this piecewise-linear approximation of the curve. For the bimanual and rational configuration space, this length is the sum of the Euclidean distance between each adjacent pair of points in the full configuration space. For $\operatorname{SO}(3)$, we use the length of the SLERP (Spherical Linear Interpolation) path since the underlying geometry is a sphere. For better results, we square the length of each segment. While this is not the same as the distance between adjacent points, the squared objective is much better numerically for the optimizer and in the limit of an infinitely-fine discretization, these will both produce the same answer [24, p.189]. A higher-resolution approximation of the cost will yield a more accurate approximation of the arc length cost, but require more computational effort. For our experiments, we find that using 10 samples strikes a good balance of accuracy and speed.
@@ III-D1 Initialization
As PGD finds local minimizers, the solution highly depends on the initialization. Within the GCS workflow, this initialization can come from two distinct candidates: after or during the rounding procedure (the step which projects the convex relaxation result to the near optimal discrete solution). Post processing the solution after rounding is a great way to quickly improve a solution to a fixed discrete path in the parametrized space. We focus on this method, but one could also use this nonconvex optimization for each sampled path as an integral component of the rounding stage.

### v2

@@ Abstract
Optimization based motion planning provides a useful modeling framework through various costs and constraints. Using Graph of Convex Sets (GCS) for trajectory optimization gives guarantees of feasibility and optimality by representing configuration space as the finite union of convex sets. Nonlinear parametrization can be used to extend this technique (to handle cases such as kinematic loops), but this often distorts distances such that convex objectives yield paths suboptimal in the original space. We present a method to extend GCS to nonconvex objectives, allowing us to “undistort” the optimization landscape while maintaining feasibility guarantees. We demonstrate our method’s efficacy on three different robotic planning domains: a bimanual robot moving an object with both arms, the set of 3D rotations using Euler angles, and a rational parametrization of kinematics that enables certifying regions as collision free. Across the board, our method significantly improves path length and trajectory duration with only a minimal increase in runtime.
Index Terms:
Motion and Path Planning, Optimization and Optimal Control, Bimanual Manipulation
@@ I Introduction
Reliable motion planning is essential to developing and deploying robotic manipulation systems. Such systems need to produce efficient paths while obeying various constraints. Optimization-based motion planning, which minimizes an objective function while satisfying constraints, offers a powerful paradigm to solve this problem. The decision variables describe the robot’s trajectory, the objective allows for choosing desired qualities in the solution, and constraints on these decision variables define obstacle avoidance, dynamic limits, and other interesting task-specific constraints such as coordinating arms in a bimanual system. However, the power of these techniques is tempered by the need to carefully formulate the optimization problems for reliability.
@@
A manipulator’s configuration space is often inherently nonconvex, and nonconvex trajectory optimization usually cannot guarantee optimality (due to local minima), or even feasibility. These guarantees are key to efficient and robust systems that can be used for repetitive motions in safety critical settings. To get a convex optimization formulation that allows for such guarantees, Graphs of Convex Sets (GCS) [1, 2] encodes the nonconvexities from obstacle avoidance as discrete decisions. Specifically, the inner approximation of planning or configuration space is represented as a series of intersecting collision free convex subsets. Then, constructing a graph (where vertices are a convex c-space set and edges connect intersecting sets) allows for searching discrete paths through these sets while simultaneously optimizing for the optimal continuous path within each set. Many relevant properties of the trajectory and its derivatives can be transcribed into convex costs and constraints [2].
@@
Figure 1: Experiments include constrained bimanual motion planning between shelves (top) and certifiable 7DoF KUKA iiwa trajectories between bins (bottom). The red path is the original result, and the blue path is our improved result.
When the configuration space does not admit finite Euclidean inner approximations, it may be possible under a change of coordinates. For example, end-effector constraints in bimanual robots such as two hands rigidly attached to the same object require planning on a nonlinear manifold in configuration space. Analytic inverse kinematics (IK) enables parameterizing convex sets on this manifold [3]. Planning over the space of 3D rotations using GCS requires a Euler angles parametrization [4, §2.7.5]. Even in cases where c-space admits inner approximations, parameterization can enable useful functionality or properties. For example, using the half tangent rational parametrization to write robot kinematics enables rigorous algebraic collision free certification [5, 6].
@@
While these parametrizations enable GCS to solve key robotics problems, they are also non-isometric; the shortest path in the parametrized space may not be the shortest path in the configuration space. This distortion leads to suboptimal results when the convex objective used in the parametrized space is a weak approximation for the true objective. Handling nonconvex objectives in GCS, such as the true objective from the original space, gives more modeling freedom and widens the breadth of problems we can tackle.
@@
This work’s key contribution is using Projected Gradient Descent to optimize nonconvex objectives in a GCS setting. A gradient based solver guarantees local optimality around the initial guess if the objective is Lipschitz-continuous. We keep constraints convex, so a small convex program can project infeasible solutions back to feasibility. By exploiting structured nonconvexity, our solver improves optimality of solutions while maintaining feasibility guarantees.
@@
For each of the three parametrizations mentioned earlier (bimanual IK, Euler angles, rational kinematics), we formulate and test a nonconvex objective against a convex surrogate in the GCS formulation. This nonconvex optimization is treated as a post-processing step, improving the best solution from GCS. Our method offers significant quantitative and qualitative improvements to motion plans across multiple experiments: path lengths and trajectory times shorten and visual artifacts of planning in the distorted parametrized space are undone. Expanding beyond just countering distortions, we also optimize over a general nonconvex cost, spatial curvature, to speed up bimanual trajectories.
@@
In the rest of the paper, we review related work on nonconvex trajectory optimization and give necessary background on GCS and nonconvexity in GCS. Then, we describe our methodology, relevant implementation details, and experimental setups. Finally, we show results, and conclude with a brief discussion of limitations of our work and potential directions for future research.
@@ II-A Graphs of Convex Sets
Graph of Convex Sets (GCS) presents a new strategy for solving the shortest path problem with continuous and discrete decisions. Formally, a GCS is a graph, where each vertex $v$ has an associated continuous variable $x_{v}$ within a convex set $X_{v}$, and each edge $(u,v)$ is a convex function of $x_{u}$ and $x_{v}$. Finding the shortest path $\mathcal{P}$ through this graph can then be formulated as a Mixed Integer Convex Problem (MICP) with $\mathcal{P}$ and $x_{v}$ as decision variables [1, §5].
GCS can be used to solve motion planning problems when given convex collision-free sets. These sets constitute the vertices of $\mathcal{P}$, and their union is an inner approximation of our planning space. The decision variable $x_{v}$ describes the continuous trajectory through the set $v$. As in GcsTrajOpt [2], we define $x_{v}$ as the control points of a Bézier curve to parametrize the continuous trajectory. This choice admits convex path continuity and differentiability constraints, and guarantees collision-avoidance of the whole trajectory [2, p.9]. GcsTrajOpt minimizes the distance between adjacent control points as a proxy for minimizing the length of the curve [1]. We use the same objective for the convex relaxation and any convex optimization we do.
The above discussion focuses only on the path, but in GcsTrajOpt, $x_{v}$ also has a time-scaling variable, $h$. We want to avoid nonconvex acceleration constraints that use this time parametrization variable. Instead, we use TOPP-RA (Time Optimal Path Parametrization based on Reachability Analysis) [18] to generate timed trajectories from our planned spatial paths. Any differentiable path can be navigated under acceleration constraints by slowing down, so using TOPP-RA maintains feasibility guarantees by keeping nonconvexity out of our constraints. So, for a collision free convex set $Q_{i}$, our vertices are of dimension $Q_{i}^{d+1}$ given Bézier curves of degree $d$ with $d+1$ control points.
@@ II-B1 IK and Constrained Bimanual Planning
Constrained bimanual manipulation, or when two robot arms move with a fixed transform between their end effectors, requires a equality constraint in task space. This non-linear inequality prevents us from using GCS on the $\mathbb{R}^{14}$ configuration manifold as it is. Cohn et al. [3] use IK to determine the joint angles of a subordinate robot given the end effector position of the leading arm. This parametrization collapses the $\mathbb{R}^{14}$ full joint space into an $\mathbb{R}^{8}$ space formed by the leading arm’s joints and a redundancy factor (required to generate consistent joint angles for the subordinate arm). However, minimizing path length in the $\mathbb{R}^{8}$ only minimizes the path length for the leading arm and ignores the subordinate arm. As a result, paths visibly favor the leading arm.
@@ II-B3 Rational Kinematics to Certify Collision Free
The forward kinematic mapping (needed for checking if a configuration is collision-free) is a trigonometric polynomial. Amice et al. [5] write this nonconvex relationship as a multi-linear polynomial, using the tangent half-angle substitution $s=\tan\frac{\theta}{2}$, further implying
@@
for $\theta\in(-\pi,\pi)$. Changing coordinates allows the formulation of Semi-Definite Programs (SDPs) to certify non-collision of regions in the robot’s rational c-space with task space obstacles. This certification can be done for individual convex sets [5], or even an entire trajectory [6].
Figure 2: A stereographic projection about $N$ projects the bottom of the black rectangle as being smaller than the top. An optimal distance planner operating in the post-projection (parametrized) space would favor the bottom despite the sides being equal in actuality. Image generated using [19].
@@ II-C Nonconvexity and GCS
There is precedent for handling nonconvexity in GCS or similar optimization frameworks. The original GcsTrajOpt preprint [20] suggested using convex approximations to incorporate nonconvex objectives and constraints. In line with this suggestion, existing GCS works using the parametrizations from Subsection II-B use a convex surrogate objective to approximate the optimal solution. While this approach preserves convexity, the approximations are inherently heuristic and must be hand-designed. Moreover, optimizing for an approximation bounds the optimality of the solution by the quality of the approximation. The nuances of the approximations can also lead to systematic pathologies, such as the imbalance between arms in the bimanual planning domain.
Another approach is to use local convex approximations of the nonconvexity. Clark and Xie [21] suggest approximating the nonconvex costs using piecewise-linear approximations and creating smaller sets within which the objective is convex. This approach maintains convexity, but may scale poorly when dealing with complex objectives and finer approximations. Using a mix of biconvex alternation and local convex approximation, Fast Path Planning [22] handles a bilinear in a similar set-up as GCS. The nonconvexity in this problem is contained to the constraints and is handled using alternation. The nonconvexity being a bilinear nonconvexity is key to enabling this method. Our work aims to enable a broader class of nonconvexity in the objective functions.
@@ III-A Nonlinear Changes of Coordinates
Each of the parametrizations from Subsection II-B distort the robot’s configuration space by introducing a nonlinear change of coordinates. More formally, each domain has a smooth (nonlinear) transformation $\alpha:Q\rightarrow C$ that maps $x$ from the more useful parametrized space $Q$ to a point $\alpha(x)$ in the original configuration space $C$. Each of the works used GCS to solve for a trajectory in $Q$, which then is remapped to $C$ using $\alpha$ to get an actual robot trajectory. However, since $\alpha$ is a nonlinear transformation, the minimum length path in $Q$ is not guaranteed to be the minimum length path in $C$.
The key limitation is that the convex path length cost in $Q$ can be arbitrarily far from the true objective: minimizing distance in $C$. Using $\alpha$ in the objective enables changing coordinates back to the original space $C$ and defining a true (now nonconvex) objective, but the sets and constraints stay convex in the parametrized space $Q$.
For the constrained bimanual case, we define $\alpha:\mathbb{R}^{8}\rightarrow\mathbb{R}^{14}$ is as the nonlinear analytic IK function with an original configuration space of both arms’ joints ($\mathbb{R}^{14}$) and a parametrized space of one arm’s joints and the self-motion of the other arm ($\mathbb{R}^{8}$). For planning over $\operatorname{SO}(3)$ with Euler angles, $\alpha:\mathbb{R}^{3}\rightarrow\mathbb{R}^{4}$ is the standard conversion from Euler angles to quaternions. For planning in the rational parametrization of kinematics, $\alpha$ is defined as $\theta=2\tan^{-1}s$.
@@ III-B Formulating the Optimization
The nonconvex objective using $\alpha$ still needs to be expressed in terms of our decision variables $x_{v}$, the control points of the Bézier curve in $Q$. We cannot directly apply $\alpha$ on $x_{v}$ to define a distance objective as the remapped control points from $Q$ do not define a same Bézier curve in $C$. However, any points along the Bézier curve in $Q$ will still be along the same path in $C$, and any point along the Bézier curve is a convex combination of its control points. Therefore, a piecewise-linear approximation of the curve in $Q$ maps to a piecewise-linear approximation in $C$ using $\alpha$.
The representative cost can then be the length of this piecewise-linear approximation. For the bimanual and rational configuration experiments, we sum the Euclidean distance between each adjacent pair of points in the full configuration space. For $\operatorname{SO}(3)$, we use the length of the Spherical Linear Interpolation (SLERP) path since the underlying geometry is a sphere. For better results, we square the length of each piece. This objective is better numerically for the optimizer and in the limit of an infinitely-fine discretization, it will both produce the same answer as a piecewise L2 norm [24, p.189]. For our experiments, using 10 samples per region to estimate the path strikes a good balance of accuracy and speed: a higher resolution approximation will be more accurate, but require more computational effort.
@@ III-D1 Initialization
As PGD finds local minimizers, the solution highly depends on the initialization. Within the GCS workflow, this initialization can come from two distinct candidates: after or during the rounding procedure (the step which projects the convex relaxation result to the near optimal discrete solution). Post-processing the solution after rounding is a great way to quickly improve a fixed discrete path in the parametrized space. We focus on this method, but one could also use this nonconvex optimization for each sampled path as an integral component of the rounding stage.